# Reliable quantum advantage in quantum battery charging

### v1

@@ Appendix B The generating function of the statistics
Similarly, we can directly prove that the term $\langle\mathscr{S}^{2}(\tau,\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\hat{a}\rangle_{\text{cav}}$ reads $N\mathscr{S}^{2}(\tau)$.
@@ Appendix C Relations between the moments and the cumulants: sketch of proof
$\begin{split}\langle\Delta U_{\tau}^{n}\rangle&=(-i)^{n}\bigl{[}ig\rho_{01}(0)(+\frac{i}{2}\hbar\omega_{\text{qub}})^{n}\langle\eta(\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\rangle_{\text{cav}}\\ &-ig\rho_{10}(0)(+\frac{i}{2}\hbar\omega_{\text{qub}})^{n}\langle\hat{a}\eta^{\ast}(\hat{a}^{\dagger}\hat{a})\rangle_{\text{cav}}\\ &+g^{2}\rho_{11}(0)(+i\hbar\omega_{\text{qub}})^{n}\langle\mathscr{S}^{2}(t,\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\hat{a}\rangle_{\text{cav}}\\ &+g^{2}\rho_{00}(0)(-i\hbar\omega_{\text{qub}})^{n}\langle\hat{a}\mathscr{S}^{2}(t,\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\rangle_{\text{cav}}\\ &-ig\rho_{01}(0)(-\frac{i}{2}\hbar\omega_{\text{qub}})^{n}\langle\eta(\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\rangle_{\text{cav}}\\ &+ig\rho_{10}(0)(-\frac{i}{2}\hbar\omega_{\text{qub}})^{n}\langle\hat{a}\eta^{\ast}(\hat{a}^{\dagger}\hat{a})\rangle_{\text{cav}}\bigr{]}\\ &\equiv\bigl{(}\hbar\omega_{\text{qub}}\bigr{)}^{n}F(\tau,n)\end{split}$ (27)

### v2

@@ Appendix B The generating function of the statistics
Similarly, we can directly prove that the term $\langle\hat{a}^{\dagger}\mathscr{S}^{2}(\tau,\hat{a}^{\dagger}\hat{a})\hat{a}\rangle_{\text{cav}}$ reads $N\mathscr{S}^{2}(\tau)$.
@@ Appendix C Relations between the moments and the cumulants: sketch of proof
$\begin{split}\langle\Delta U_{\tau}^{n}\rangle&=(-i)^{n}\bigl{[}ig\rho_{01}(0)(+\frac{i}{2}\hbar\omega_{\text{qub}})^{n}\langle\hat{a}^{\dagger}\eta(\hat{a}^{\dagger}\hat{a}+\hat{\mathds{I}})\rangle_{\text{cav}}\\ &-ig\rho_{10}(0)(+\frac{i}{2}\hbar\omega_{\text{qub}})^{n}\langle\eta^{\ast}(\hat{a}^{\dagger}\hat{a}+\hat{\mathds{I}})\hat{a}\rangle_{\text{cav}}\\ &+g^{2}\rho_{11}(0)(+i\hbar\omega_{\text{qub}})^{n}\langle\hat{a}^{\dagger}\mathscr{S}^{2}(t,\hat{a}^{\dagger}\hat{a}+\hat{\mathds{I}})\hat{a}\rangle_{\text{cav}}\\ &+g^{2}\rho_{00}(0)(-i\hbar\omega_{\text{qub}})^{n}\langle\hat{a}\mathscr{S}^{2}(t,\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\rangle_{\text{cav}}\\ &-ig\rho_{01}(0)(-\frac{i}{2}\hbar\omega_{\text{qub}})^{n}\langle\eta(\hat{a}^{\dagger}\hat{a})\hat{a}^{\dagger}\rangle_{\text{cav}}\\ &+ig\rho_{10}(0)(-\frac{i}{2}\hbar\omega_{\text{qub}})^{n}\langle\hat{a}\eta^{\ast}(\hat{a}^{\dagger}\hat{a})\rangle_{\text{cav}}\bigr{]}\\ &\equiv\bigl{(}\hbar\omega_{\text{qub}}\bigr{)}^{n}F(\tau,n)\end{split}$ (27)